# VidGuard-R1: AI-Generated Video Detection and Explanation via Reasoning MLLMs and RL

### v2

@@ 4.3 Case studies on explanations
For instance, in Figure 4, the smooth hand motion initially suggests realism; however, once the origami folds in a physically implausible manner, the model revises its judgment. In Figure 4, although the lighting and shadows are consistent—typically a cue for authenticity—the model recognizes that this is insufficient in a largely static scene with only a stationary machine and a human hand. In particular, even in its final prediction, the model reflects on earlier realistic cues and acknowledges that the video could still be real, underscoring its nuanced, human-like reasoning in assessing video quality. Additional case studies are provided in the Appendix D.
@@ 5.1 Limitations
Our dataset currently includes fake videos generated exclusively with the HunyuanVideo and CogVideoX. Although it is constructed in a pairwise fashion to ensure contextual similarity between real and fake videos, its generalizability remains limited. Expanding the dataset to incorporate fake videos from a broader range of generative models would result in a more diverse and robust training set, thereby improving its applicability to real-world scenarios.
@@ Appendix B Comprehensive Benchmark Evaluation
F3Net (Qian et al., 2020a) CNN 37.43 37.27 36.46 39.59 52.76 42.52
I3D (Carreira & Zisserman, 2017) CNN 8.15 8.29 60.11 59.24 93.99 49.23
@@
VidGuard-R1 (GRPO,
GenVideo-pretrained, Zero-shot)
MLLM 97.24 96.59 99.88 99.93 88.14 96.37
@@
VidGuard-R1 (GRPO,
GenVidBench-pretrained, Zero-shot)
MLLM R 0.95 0.98 0.90 0.89 0.97 0.85 0.99 0.93 0.81 0.87 0.92
@@
Appendix C Prompt
@@
Appendix D Case studies on explanations
D.1 GenVidBench
@@
D.2 GenVideo
Figures 1322 show inference examples for videos generated by ten different AI models included in the GenVideo testing dataset: Sora (OpenAI, 2024), Morph Studio (mor, 2025), Gen2 (Esser et al., 2023a), HotShot (hot, 2025), Lavie (Wang et al., 2025b), Show-1 (Zhang et al., 2024a), Moonvalley (moo, 2025), Crafter (Chen et al., 2023), ModelScope (Wang et al., 2023a), and DreamVideo (Wei et al., 2024).

### v3

@@ 4.4 Case studies on explanations
For instance, in Figure 4, the smooth hand motion initially suggests realism; however, once the origami folds in a physically implausible manner, the model revises its judgment. In Figure 4, although the lighting and shadows are consistent—typically a cue for authenticity—the model recognizes that this is insufficient in a largely static scene with only a stationary machine and a human hand. In particular, even in its final prediction, the model reflects on earlier realistic cues and acknowledges that the video could still be real, underscoring its nuanced, human-like reasoning in assessing video quality. Additional case studies are provided in the Appendix E.
@@ 5.1 Limitations
Our dataset currently contains fake videos generated by HunyuanVideo and CogVideoX, the main open source models that support large-scale text-image joint conditioning. Most other diffusion models offer only text or image conditioning, making them less suitable for our pairwise data construction. Although this design ensures strong contextual alignment, adding outputs from more generative models would increase diversity and robustness, improving real-world applicability.
Ethics Statement
This work does not involve personally identifiable information or sensitive user data. All datasets used in our experiments are publicly available and constructed in accordance with their licenses and usage guidelines. VidGuard-R1 is designed to mitigate societal risks associated with AI-generated videos, such as misinformation and reputational harm, by providing interpretable CoT reasoning to assist human verification. To the best of our knowledge, the method does not introduce risks related to fairness, safety, or privacy.
Acknowledgement
This work was supported in part by the National Science Foundation under Grant CNS-2212297. We are grateful to the anonymous reviewers for their insightful suggestions, which helped improve the quality and clarity of this paper.
@@ Appendix B Comprehensive Benchmark Evaluation
F3Net (Qian et al., 2020) CNN 37.43 37.27 36.46 39.59 52.76 42.52
I3D (Carreira and Zisserman, 2017) CNN 8.15 8.29 60.11 59.24 93.99 49.23
@@
VidGuard-R1 (GRPO, GenVideo-pretrained, Zero-shot) MLLM 97.24 96.59 99.88 99.93 88.14 96.37
@@
VidGuard-R1 (GRPO, GenVidBench-pretrained, Zero-shot) MLLM R 0.95 0.98 0.90 0.89 0.97 0.85 0.99 0.93 0.81 0.87 0.92
@@
Appendix C Zero-Shot Generalization to Unseen Generative Models
To evaluate the robustness of VidGuard-R1 beyond the curated training sources, we assess its zero-shot performance on a diverse set of recently released generative video models that were not used during training, including Gen-3 Alpha (19), Pika (42), Pika 2.2 (42), Luma Ray2 (32), Sora (39), Veo2 (20), Veo3 (20), and Wan 2.1 (48). Table 12 summarizes performance across these unseen systems. VidGuard-R1 achieves accuracy above 80% in all cases, reaching up to 96.36%, demonstrating strong generalization to more recent and increasingly realistic generative models.
Table 12: Zero-shot detection accuracy on unseen generative models
Model Total Correct Incorrect Accuracy (%)
Gen-3 Alpha 56 49 7 87.50
Pika 110 101 9 91.82
Pika 2.2 110 106 4 96.36
Luma Ray2 110 98 12 89.09
Sora 110 102 8 92.73
Veo2 52 45 7 86.54
Veo3 55 45 10 81.82
Wan2.1 55 46 9 83.64
These analyses demonstrate that VidGuard-R1 generalizes effectively beyond the curated generative sources and remains robust across a wide range of unseen, high-quality video generation models.
Appendix D Prompt
@@
Appendix E Case studies on explanations
E.1 GenVidBench
@@
E.2 GenVideo
Figures 1322 show inference examples for videos generated by ten different AI models included in the GenVideo testing dataset: Sora (39), Morph Studio (35), Gen2 (Esser et al., 2023), HotShot (24), Lavie (Wang et al., 2025), Show-1 (Zhang et al., 2024a), Moonvalley (34), Crafter (Chen et al., 2023), ModelScope (Wang et al., 2023a), and DreamVideo (Wei et al., 2024).